OpenStudy (melodious):

The moon's mass is approximately 1% of the earth's mass. How does the gravitational pull of the earth on the moon compare with the gravitational pull of the moon on the earth?

OpenStudy (anonymous):

Let Mass of Earth be Me Mass of Moon Mm Gravitational Pull of Earth Fe and Gravitational Pull of Moon Fm \[M_{e},M_{m},F_{e},F_{m}\] These are the quantities Now the gravitational pull is proportional to the mass of the body so \[F \propto M\] Removing the proportionality we introduce a constant k \[F=kM\] Now for earth and moon we have the equations \[F_{e}=kM_{e}\] \[F_{m}=kM_{m}\] Dividing these we get \[\frac{F_{e}}{F_{m}}=\frac{M_{e}}{M_{m}}\] From this we get \[F_{m}=\frac{M_{m}}{M_{e}}.F_{e}\] But you are also given Mm as 1% of Me so \[M_{m}=\frac{1}{100}.M_{e} \implies M_{e}=100M_{m}\] Substitute this value in equation for Fm and simplify

OpenStudy (melodious):

can explain this to me pls @Nishant_Garg